# Subhalos are Distributed Anisotropically About Their Hosts

### v1

@@ 5.1 Principle Axis and Angular Momentum Alignments of Subhalos and Their Hosts
In principle, it may be possible for the major axes of host halo and subhalo components to be aligned without a strongly anisotropic distribution of subhalos (though this seems inconsistent with our other results, see below). Suppose that the distribution of subhalos is characterized by principle axis ratios $(b/a)_{\mathrm{subs}}$ and $(c/a)_{\mathrm{subs}}$ (from the subhalo only samples) that are significantly larger than the principle axis ratios determined by the overall mass of the host halos (from the host only) samples. For example, if $(c/a)_{\mathrm{subs}}\sim 1\gg(c/a)$ and $(b/a)_{\mathrm{subs}}\sim 1\gg(b/a)$, then the subhalo mass would be distributed nearly isotropically, so that axis alignments would be far less noteworthy. However, this is not the case. We find that subhalo mass is distributed approximately as anisotropically as its host halo’s mass as quantified by axis ratios. The axis ratios for host halos in the Milky Way-mass simulations are $(b/a)=0.86\pm 0.07$ and $(c/a)=0.76\pm 0.07$, while the corresponding ratios for subhalo mass are $(b/a)_{\mathrm{subs}}=0.8\pm 0.1$ and $(c/a)_{\mathrm{subs}}=0.7\pm 0.1$. Similarly, for cluster mass halos the axis ratios are $(b/a)=0.78\pm 0.09$ and $(c/a)=0.68\pm 0.07$ for host halos while they are $(b/a)_{\mathrm{subs}}=0.8\pm 0.1$ and $(c/a)_{\mathrm{subs}}=0.6\pm 0.1$ for the subhalos. Subhalo mass is distributed at least as anisotropically as the host halo mass indicating that alignments among principle axes designate meaningful alignments of a strongly anisotropic distribution of subhalos.
@@
Figure 3: The cumulative fraction of halos with angle <$|\rm cos(\theta)|$ between the host halo major axis $\vec{A}_{\mathrm{Host}}$ and that of its subhalos, $\vec{A}_{\mathrm{Sub}}$, as a function of $|\rm cos(\theta)|$. The solid line is the result from the simulation data, the dashed line and the filled area are the CDF for an isotropic distribution and the area within the 14th and 86th percentiles. Panel (a) is for Milky Way mass halos and (b) is for cluster mass halos. The Kolmogorov-Smirnov (KS) probability that the solid line distribution in (a) is selected from an isotropic distribution is $P_{\rm KS}$ is 6.38 $\times$ 10-15. For panel (b) this probability is 1.88 $\times$ 10-23.
@@
In Fig. 4 we compare the angular momentum vector of the host halo to that of its subhalos. Additionally, the x-axis in Fig. 4 is $\rm cos(\theta)$ rather than $|\rm cos(\theta)|$ to account for cases in which subhalo angular momenta are anti-parallel to that of their host. The top panel corresponds to Milky Way-mass halos and the bottom to cluster-mass halos. We see a similar trend as in Fig. 3 in which there is a deficit of halos at smaller values of $\rm cos(\theta)$, indicating that there is alignment between $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$. The KS probability that the distribution for Milky Way-mass halos in 4(a) is selected from an isotropic distribution is $P_{\rm KS}$ is $\sim 1\times 10^{-\rm 9}$. For cluster- mass halos in 4(b) $P_{\rm KS}$ is $\sim 4\times 10^{-\rm 5}$. As we discussed in the preceding paragraph, the tendency for subhalo and host halo angular momenta to be aligned makes sense within the context of hierarchical growth – subhalos merging in from coherent directions along filaments will preferentially orbit near a plane containing the major axis. The fact that we detect alignment between major axes strengthens this argument.
@@
Finally, we quantify the alignment between angular momenta with the median angular separation between host and subhalo angular momentum vectors. For Milky Way-mass halos, the median angular separation is $\theta$ = $43\pm 7.6^{\circ}$ or cos($\theta$) = 0.73 $\pm$ 0.09. For Cluster-mass halos, it is $\theta$ = $56\pm 5.7^{\circ}$ or cos($\theta$) = 0.56 $\pm$ 0.08. Errors are calculated by bootstrapping using the same method as in Section 4. For an isotropic distribution of angles (spanning from 0 to 180), we would expect a median of 90, with 0 indicative of perfect alignment between the angular momentum vectors, and 180 indicative of $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$ pointing in complete opposite directions. Our findings show that the median angle of separation between angular momentum of the host and subhalos is significantly smaller than 90, indicative of strong alignment between the two, and is consistent with the discussion in the previous paragraph that subhalos should orbit in the same plane as the host halo.
Figure 4: The same as Fig. 3 except for the angles between the host and subhalo angular momentum vectors, $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$, respectively. The Kolmogorov-Smirnov (KS) probability that the distribution in panel (a) is selected from an isotropic distribution is $P_{\rm KS}$ is 1.29 $\times$ 10-9. The probability that the distribution in panel (b) is selected from an isotropic distribution is $P_{\rm KS}$ is 0.0028.
@@
Although we anticipate that the alignments of $\vec{A}_{\mathrm{Sub}}$ and $\vec{A}_{\mathrm{Host}}$, and, $\vec{J}_{\mathrm{Sub}}$ and $\vec{J}_{\mathrm{Host}}$ are a result of mergers along coherent directions, such as filaments, we should expect significant scatter in these alignments. Cluster-sized halos are comparable in size to the filament in which they reside and may also be located at a node connecting multiple filaments. Galaxy-sized halos at present are often smaller than the size of the filaments in which they live but may have accreted their surviving subhalos at times when filaments were thinner (Vera-Ciro et al., 2011). The distribution of subhalos that merge from other directions or from presently-subdominant filamentary structure introduces considerable scatter into the anisotropic distributions we have found. Furthermore, simulations have shown that spin alignment with filaments has a mass dependence (Bailin & Steinmetz, 2005; Aragón-Calvo et al., 2007; Paz et al., 2008; Zhang et al., 2009; Codis et al., 2012; Libeskind et al., 2011; Libeskind et al., 2015; Forero-Romero et al., 2014). In particular, Codis et al. (2012) find that the orientation of a halo’s spin relative to the filament in which it resides has a mass dependence with a turnover from pointing parallel to the filament to perpendicular at a critical mass of $\sim$ 5($\pm$1) $\times$ $10^{12}\text{M}_{\odot}$. This mass is greater than that of our Milky way-mass halos and considerably lower than that of our cluster-mass halos. This suggests that Milky Way-mass halos will be oriented differently relative to the present-day large-scale structure than cluster-mass halos will. This is interesting to note and may be important for a variety of forthcoming applications, including gravitational lensing studies, but we will not pursue this further in this work.
@@ 5.2 The Distribution of Subhalo Mass as a Function of Viewing Angle
In order to interpret strong gravitational lensing measurements, it is most practical to focus on the mass in the central region of the halo and specifically, the mass within approximately the Einstein radius in projection. Because of the ellipsoidal shape of halos, the projected matter density near the center will vary depending on the orientation from which the halo is viewed and a greater projected matter density will be observed near the center when the halo is viewed along its long axis compared to its shortest. We predict that a greater projected density of subhalos at the halo center will also be observed when the halo is viewed along this axis.
@@
Determining the influence of subhalos on strong lensing systems is a topic of great contemporary interest. We have already demonstrated that subhalos are distributed anisotropically around their hosts and that subhalos are found preferentially aligned the long axes of their hosts. We quantify this effect to inform its importance in previous and forthcoming lensing measurements. We show the results from this analysis in Fig. 5 where the top row correspond to Milky Way-mass halos and the bottom row to cluster-mass halos. The panels of Fig. 5 depict the mass fraction in subhalos in a cylindrical projection viewed from an angle $\theta$ relative to the major axis of the halo. The left-hand panels show a cylinder radius of $10\%$ of the host halo virial radius, while the right-hand panels show results for a cylinder radius of $5\%$ of the host halo virial radius. The horizontal axis is the absolute value of the cosine of the angular separation between the cylinder axis and the halo major axis, |$\rm{cos}(\theta)$|. The cylinder scheme was discussed in Section 3 and it may be useful to review Fig. 1 to visualize the geometry under consideration. In all panels, the vertical axis is the ratio of the projected subhalo mass fraction enclosed within a particular cylinder, oriented at an angle $\theta$ away from the halo major axis, divided by the median mass fraction for all cylinders across all halos and angles.
@@
Figure 5: The four panels show subhalo mass fraction captured within cylinders projected along axes randomly oriented to the major axis normalized by the median mass fraction across all halos and angles. The top two panels in blue correspond to Milky Way-mass halos and the two bottom pink ones to cluster-mass halos. The x-axis is the absolute value of the cosine of the angular separation between the cylinder axis and the halo major axis, and, the y-axis is the normalized subhalo mass fraction. The solid line is the median normalized mass fraction within bins of $|\cos(\theta)|$. The shaded regions are the error margins corresponding to 14th and 86th percentiles.
@@
In each of the panels of Fig. 5, the solid line is the median normalized subhalo mass fraction at viewing angle |cos($\theta$)| and the shaded region is the area corresponding to the error on the median. Mass fractions are normalized by the median subhalo mass fraction across all halos and angles. Consequently, if the subhalos were distributed isotropically, the solid lines in each panel would be horizontal lines with a value of 1.0. In both Milky Way-mass and cluster-mass halos, the subhalo mass fraction increases as cylinder orientation approaches the major axis. This effect is more dramatic for cluster-mass halos, whereas the trend for Milky Way-mass halos is both noisier (due to the smaller sample size) and smaller. This result naturally emerges from the anisotropies already discussed in Section 5.1. We saw overall stronger alignments with particular axes for cluster-mass halos in Section 5.1 as well, so the difference in the strength of this effect between Milky Way-mass halos and cluster-mass halos is not surprising. We attribute the noisiness in the Milky Way result to the fact that there are fewer than half as many Milky Way-mass host halos compared to cluster-mass host halos. Moreover, smaller host halos tend to have a smaller total fraction of their mass in subhalos, so the Milky Way-mass hosts contain fewer subhalos and thus there is a larger shot noise associated with subhalo position.
@@
For lensing, a massive subhalo of some mass, $M$, in the line of sight will leave a larger lensing signal than a collection of smaller subhalos with collective mass, $M$. To this end, we repeated the analysis discussed above with a variety of subsamples computed using threshold samples of subhalos above a minimum mass. We calculate projected subhalo mass for subhalos of mass, $M$, greater than 0.025, 0.05, 0.075, and 0.1 times the total halo virial mass. These numbers are selected such that they fall within the range of available subhalo masses – Milky Way-mass halos in our sample do not have subhalos more massive than 0.2 times their total halo virial mass and only a handful of cluster-mass halos have subhalos that exceed this threshold. We find that as this lower mass limit is increased, there is a small increase in the overabundance of projected subhalo mass near the major axis, although, the change is not statistically significant. When considered together with our other results indicating subhalo anisotropies, we nevertheless encourage the reader to be mindful of this potential effect.

### v2

@@ 5.1 Principle Axis and Angular Momentum Alignments of Subhalos and Their Hosts
In principle, it may be possible for the major axes of host halo and subhalo components to be aligned without an anisotropic distribution of subhalos (though this seems inconsistent with our other results, see below). Suppose that the distribution of subhalos is characterized by principle axis ratios $(b/a)_{\mathrm{subs}}$ and $(c/a)_{\mathrm{subs}}$ (from the subhalo only samples) that are significantly larger than the principle axis ratios determined by the overall mass of the host halos (from the host only) samples. For example, if $(c/a)_{\mathrm{subs}}\sim 1\gg(c/a)$ and $(b/a)_{\mathrm{subs}}\sim 1\gg(b/a)$, then the subhalo mass would be distributed nearly isotropically, so that axis alignments would be far less noteworthy. However, this is not the case. We find that subhalo mass is distributed approximately as anisotropically as its host halo’s mass as quantified by axis ratios. The axis ratios for host halos in the Milky Way-mass simulations are $(b/a)=0.86\pm 0.10$ and $(c/a)=0.74\pm 0.11$, while the corresponding ratios for subhalo mass are $(b/a)_{\mathrm{subs}}=0.84\pm 0.16$ and $(c/a)_{\mathrm{subs}}=0.66\pm 0.16$. Similarly, for cluster mass halos the axis ratios are $(b/a)=0.78\pm 0.08$ and $(c/a)=0.68\pm 0.07$ for host halos while they are $(b/a)_{\mathrm{subs}}=0.60\pm 0.2$ and $(c/a)_{\mathrm{subs}}=0.36\pm 0.18$ for the subhalos. Subhalo mass is distributed at least as anisotropically as the host halo mass indicating that alignments among principle axes designate meaningful alignments of a strongly anisotropic distribution of subhalos. In fact, our results suggest that the distribution of subhalo mass is somewhat less spherical than the overall mass distribution of a halo.
@@
Figure 4: The cumulative fraction of halos with angle $|\rm cos(\theta)|$ between the host halo major axis $\vec{A}_{\mathrm{Host}}$ and that of its subhalos, $\vec{A}_{\mathrm{Sub}}$, as a function of $|\rm cos(\theta)|$. The solid line is the result from the simulation data, the dashed line and the filled area are the CDF for an isotropic distribution and the area within the 14th and 86th percentiles. Panel (a) is for Milky Way mass halos and (b) is for cluster mass halos.
@@
In Fig. 5, we compare the angular momentum vectors of host halos, $\vec{J}_{\mathrm{Host}}$, to that of their subhalos, $\vec{J}_{\mathrm{Sub}}$. $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$ are both defined in Section 3.2. Notice that, unlike orientations relative to the major axis, the x-axis in Fig. 5 is $\rm cos(\theta)$ rather than $|\rm cos(\theta)|$ to display cases in which subhalo angular momenta are anti-parallel to that of their host. The left panels correspond to Milky Way-mass halos and the right to cluster-mass halos. In the top row, figures $a$ and $b$, we see a similar trend as in Fig. 4 in which there is a deficit of halos at smaller values of $\rm cos(\theta)$, indicating that there is alignment between $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$. The KS probability that the distribution for Milky Way-mass halos in the top panel of 5(a) is selected from an isotropic distribution is $P_{\rm KS}$ $\sim 1\times 10^{-\rm 9}$. For cluster-mass halos in the top panel of 5(b), $P_{\rm KS}$ is $\sim 4\times 10^{-\rm 5}$. As we discussed in the preceding paragraph, the tendency for subhalo and host halo angular momenta to be aligned makes sense within the context of hierarchical growth – subhalos merging in from coherent directions along filaments will preferentially orbit near a plane containing the major axis. The fact that we detect alignment between major axes strengthens this argument.
The other two rows of panels in Fig. 5 show our results for alternative methods of measuring the subhalo angular momentum. In the first case, we define the angular momentum of each subhalo using the position and velocity of its center as specified in the halo catalog and sum over that angular momentum for all subhalos within a host to find the total angular momenta of the subhalo component of the halo. We will refer to this measure of angular momentum as $\vec{J}_{\mathrm{peak}}$. This method is distinct from the previous method in which the subhalo angular momenta were computed by summing over the particle content of each subhalo. In panels $c$ and $d$ we show the cumulative distribution of angles between $\vec{J}_{\mathrm{peak}}$ and $\vec{J}_{\mathrm{Host}}$. The distributions in $c$ and $d$ are, unsurprisingly, both similar to those of $a$ and $b$. This suggests that our alignment results are insensitive to the details of the method used to determine subhalo angular momenta. In the second alternative case, we display the alignments of angular momenta for individual subhalos, rather than for populations of subhalos associated with individual hosts. In particular, we compute the angle between the angular momentum of each individual subhalo and the angular momentum vector of its host. To distinguish this measurement, we refer to the angular momentum of a single subhalo as $\vec{J}_{\mathrm{ind}}$. We sum over the angular momenta of all particles in any individual subhalo to determine $\vec{J}_{\mathrm{ind}}$. The result of this exercise is shown in the bottom panels, $e$ and $f$. The resulting distribution of alignments differs from those in panels $a$, $b$, $c$, and $d$, because in this case the individual subhalos corresponding to a single host can fall into distinct bins of zenith angle, whereas in the previous calculation each host had a single zenith angle corresponding to the alignment of the host angular momentum with the net angular momentum of the population of subhalos. This has the effect of spreading the contribution from the subhalo population associated with any particular host galaxy among a range of zenith angle values. Furthermore, the alignment distribution computed in this way is significantly less noisy because the distribution is assembled from many more data points. Calculating angles for individual halos allows us to decipher whether or not the alignment in the top two figures, $a$ and $b$, is due to a large fraction of subhalos having angular momenta aligned with that of their hosts, or, subhalo angular momenta that are spread out but average to a value close to that of its host. The anisotropy is weaker for $\vec{J}_{\mathrm{ind}}$, implying that those subhalos with the largest angular momenta are more well-aligned with the host angular momentum.
@@
Finally, we quantify the alignment between angular momenta with the median angular separation between host and subhalo angular momentum vectors. For Milky Way-mass halos, the median angular separation is $\theta$ = $50\pm 9^{\circ}(38.0^{\circ})$ or cos($\theta$) = 0.63 $\pm$ 0.12 (0.55). For Cluster-mass halos, it is $\theta$ = $56\pm 5.5^{\circ}(34^{\circ})$ or cos($\theta$) = 0.56 $\pm$ 0.08 (0.49). Errors are calculated by bootstrapping using the same method as in Section 4. For an isotropic distribution of angles (spanning from 0 to 180), we would expect a median of 90, with 0 indicative of perfect alignment between the angular momentum vectors, and 180 indicative of $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$ pointing in complete opposite directions. Our findings show that the median angle of separation between angular momentum of the host and subhalos is significantly smaller than 90, indicative of strong alignment between the two, and is consistent with the discussion in the previous paragraph that subhalos should orbit in the same plane as the host halo.
Figure 5: The same as Fig. 4 except for the angles between the host and subhalo angular momentum vectors, $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$, respectively. In the bottom 4 figures, we use alternative measurements of subhalo angular momentum. For comparison with the original method of measuring subhalo angular momentum, we also plot the lines from the top plots in a lighter shade along side the alternative method result. In the middle row, subhalo angular momenta is defined by the position and velocity of its density peak. Like in the top row figures, we sum over all subhalo angular momenta to define $\vec{J}_{\mathrm{peak}}$. In the bottom row, $\vec{J}_{\mathrm{ind}}$ corresponds to the angular momentum of a single subhalo found by summing over all its particles rather than the sum of particles over all subhalos belonging to a host.
@@
Although we anticipate that the alignments of $\vec{A}_{\mathrm{Sub}}$ and $\vec{A}_{\mathrm{Host}}$, and, $\vec{J}_{\mathrm{Sub}}$ and $\vec{J}_{\mathrm{Host}}$ are a result of mergers along coherent directions, such as filaments, we should expect significant scatter in these alignments. Cluster-sized halos are comparable in size to the filament in which they reside and may also be located at a node connecting multiple filaments. Galaxy-sized halos at present are often smaller than the size of the filaments in which they live but may have accreted their surviving subhalos at times when filaments were thinner (Vera-Ciro et al., 2011). The distribution of subhalos that merge from other directions or from presently-subdominant filamentary structure introduces considerable scatter into the anisotropic distributions we have found. Furthermore, simulations have shown that spin alignment with filaments has a mass dependence (Bailin & Steinmetz, 2005; Aragón-Calvo et al., 2007; Paz et al., 2008; Zhang et al., 2009; Codis et al., 2012; Libeskind et al., 2011, 2015; Forero-Romero et al., 2014). In particular, Codis et al. (2012) find that the orientation of a halo’s spin relative to the filament in which it resides has a mass dependence with a turnover from pointing parallel to the filament to perpendicular at a critical mass of $\sim$ 5($\pm$1) $\times$ $10^{12}\text{M}_{\odot}$. This mass is greater than that of our Milky way-mass halos and considerably lower than that of our cluster-mass halos. This suggests that Milky Way-mass halos will be oriented differently relative to the present-day large-scale structure than cluster-mass halos will. This is interesting to note and may be important for a variety of forthcoming applications, including gravitational lensing studies, but we will not pursue this further in this work.
@@ 5.2 The Distribution of Subhalo Mass as a Function of Viewing Angle
In order to interpret strong gravitational lensing measurements, it is most practical to focus on the mass in the central region of the halo and specifically, the mass within approximately the Einstein radius in projection. For lenses at z=0.5 composed of a galaxy within a dark matter halo within the mass range of our Milky Way-mass halos, with a source at z=2, we estimate an Einstein radius of approximately 2 arcsecs or 12.2 kpc and for lenses composed of a galaxy in a cluster-mass dark matter halo they are approximately 10 arcsecs or 61 kpc. Because of the ellipsoidal shape of halos, the projected matter density near the center will vary depending on the orientation from which the halo is viewed and a greater projected matter density will be observed near the center when the halo is viewed along its long axis compared to its shortest. We predict that a greater projected density of subhalos at the halo center will also be observed when the halo is viewed along this axis.
@@
Determining the influence of subhalos on strong lensing systems is a topic of great contemporary interest. We have already demonstrated that subhalos are distributed anisotropically around their hosts and that subhalos are found preferentially aligned the long axes of their hosts. We quantify this effect to inform its importance in previous and forthcoming lensing measurements. We show the results from this analysis in Fig. 6 where the left panels correspond to Milky Way-mass halos and the right panels to cluster-mass halos. The panels of Fig. 6 depict the mass fraction in subhalos in a cylindrical projection as a function of both radius of the cylindrical projection (top panels) and viewing angle relative to the major axis of the halo (bottom panels). The cylinder scheme was discussed in Section 3 and it may be useful to review Fig. 2 to visualize the geometry under consideration. In all panels, the vertical axis is the ratio of the projected subhalo mass fraction enclosed within a particular cylinder, oriented at an angle $\theta$ away from the halo major axis, divided by the median mass fraction for all cylindrical projections across all halos and angles.
@@
Figure 6: The top two panels show subhalo mass fraction captured within cylinders projected along predefined axes on a uniformly spaced grid on the surface of a sphere normalized by the median mass fraction across all halos and angles. The left panel in blue corresponds to Milky Way-mass halos and the right panel in pink to cluster-mass halos. The x-axis is the ratio of the radius of the cylinder to virial radius of the halo, and, the y-axis is the normalized subhalo mass fraction. The solid lines are the median normalized mass fraction within bins of $r/r_{\rm{VIR}}$ at four different angular separations ranging from aligned with the major axis (lightest shaded line) to perpendicular to the major axis (darkest shaded line). The bottom two panels contain the same information presented in a different way. For these figures, the x-axis is the angular separation. The solid and dotted lines correspond to the mass fraction within cylinders of radii equal to approximately 10% and 5% of the virial radius. The shaded regions are the error margins corresponding to 14th and 86th percentiles.
In the top two panels of Fig. 6, the solid line is the median normalized subhalo mass fraction within bins of $r/r_{\rm{VIR}}$ at four different angular separations ranging from aligned with the major axis (lightest shaded line) to perpendicular to the major axis (darkest shaded line). The shaded regions are the error margins corresponding to 14th and 86th percentiles. Mass fractions are normalized by the median subhalo mass fraction across all halos and angles. Consequently, if the subhalos were distributed isotropically, the solid lines in each panel would be horizontal lines with a value of 1.0. In both Milky Way-mass and cluster-mass halos, the subhalo mass fraction increases as cylinder orientation approaches the major axis. Smaller cylinder radii lead to dramatically larger differences, which makes sense because smaller cylinder radii correspond to averaging over a smaller physical scale so that correlations are not washed out by subhalos that are poorly aligned with the major axis, but fall within the cylinder radius. This effect is more dramatic for cluster-mass halos where the mass fraction is approximately 1.95 times greater at small cylinder radii along the major axis compared to perpendicular, whereas the trend for Milky Way-mass halos is both smaller (1.75 times greater at small radii along the major axis) and noisier (due to the smaller sample size). The bottom two panels show the median normalized mass fraction within bins of separation angle at two cylinder radii equal to approximately 5% and 10% of the virial radius. Once again, the shaded regions are the bootstrapped error margins corresponding to 14th and 86th percentiles. In these figures we again note an increase in mass fraction as cylinder orientation approaches the major axis for the same reason as discussed above. We note that the difference in maximum mass fraction between the top and bottom figures is due to the fact that in the bottom figures we are binning data in cos($\theta$).
@@
This result naturally emerges from the anisotropies already discussed in Section 5.1. We saw overall stronger alignments with particular axes for cluster-mass halos in Section 5.1 as well, so the difference in the strength of this effect between Milky Way-mass halos and cluster-mass halos is not surprising. We attribute the noisiness in the Milky Way result to the fact that there are fewer than half as many Milky Way-mass host halos compared to cluster-mass host halos. Moreover, smaller host halos tend to have a smaller total fraction of their mass in subhalos, so the Milky Way-mass hosts contain fewer subhalos and thus there is a larger shot noise associated with subhalo position.
@@
For lensing, a massive subhalo of some mass, $M$, in the line of sight will leave a larger lensing signal than a collection of smaller subhalos with collective mass, $M$. To this end, we ran a smaller scale version of the analysis discussed above with a variety of subsamples computed using threshold samples of subhalos above a minimum mass. We calculate projected subhalo mass for subhalos of mass, $M$, greater than 0.025, 0.05, 0.075, and 0.1 times the total halo virial mass. These numbers are selected such that they fall within the range of available subhalo masses – Milky Way-mass halos in our sample do not have subhalos more massive than 0.2 times their total halo virial mass and only a handful of cluster-mass halos have subhalos that exceed this threshold. For this iteration of the analysis, we calculate the mass fraction at only two radii on the scale of the expected Einstein radius, 0.05 and 0.1 times the halo virial radius, along 500 randomly selected axes through the halo. We find that as this lower mass limit is increased, there is a small increase in the overabundance of projected subhalo mass near the major axis, although, the change is not statistically significant. When considered together with our other results indicating subhalo anisotropies, we nevertheless encourage the reader to be mindful of this potential effect.